# Haar-Laplacian for directed graphs11footnotemark: 1

### v1

@@ Abstract
This paper introduces a novel Laplacian matrix aiming to enable the construction of spectral convolutional networks and to extend the signal processing applications for directed graphs. Our proposal is inspired by a Haar-like transformation and produces a Hermitian matrix which is not only in one-to-one relation with the adjacency matrix, preserving both direction and weight information, but also enjoys desirable additional properties like scaling robustness, sensitivity, continuity, and directionality. We take a theoretical standpoint and support the conformity of our approach with the spectral graph theory. Then, we address two use-cases: graph learning (by introducing HaarNet, a spectral graph convolutional network built with our Haar-Laplacian) and graph signal processing. We show that our approach gives better results in applications like weight prediction and denoising on directed graphs.
@@
directed graphs, Laplacian
@@ 1 Introduction
We assume that the graph has no self loops, hence $a_{uu}=0$, $\forall u\in\mathcal{V}$. We also assume that the graph is connected; there are no isolated vertices.
@@
where $|A|$ is the matrix whose elements are $|a_{uv}|$ and $\bm{1}$ is a vector with all elements equal to 1. When all weights are equal to one, this is the degree matrix.
The Laplacian associated to the graph is
$L=D-A$ (3)
@@
If the graph is undirected with nonnegative weights, the Laplacian enjoys nice spectral properties, which proved useful in graph signal processing [1, 2, 3]. In particular, both Laplacians are symmetric and positive semidefinite; the eigenvalues of the normalized Laplacian (4) lie in the interval $[0,2]$. The case of negative weights in undirected graphs is covered in [4].
Contribution. We propose here a new construction, called Haar-Laplacian, based on the symmetrized and skew-symmetrized adjacency matrices associated with a directed graph. Like other Laplacians for directed graphs [5, 6, 7], the Haar-Laplacian is a Hermitian matrix and thus is has real eigenvalues and orthogonal eigenvectors. We show that the Haar-Laplacian enjoys some properties that former Laplacians do not share. Moreover, we use the Haar-Laplacian in two types of applications. For graph learning, we introduce a spectral graph convolutional network called HaarNet and show its behavior for several instances of the link prediction problem in directed graphs; the best results are obtained for weight prediction, where the one-to-one relation of the Haar-Laplacian with the adjacency matrix gives it an edge over the existing approaches that do not have this property. We also illustrate the benefits of the Haar-Laplacian in graph signal processing tasks like denoising. A related work is [8], where we have used the symmetrized and skew-symmetrized adjacency matrices for building separate blocks in an autoencoder network dedicated to anomaly detection; no connection with the Laplacian was made; no results, theoretical or experimental, are replicated in this paper.
Contents. In Sec. 2, we review the state-of-the-art in the topic of Laplacians for directed graphs, from two perspectives, those of graph learning and graph signal processing. In Sec. 3, we introduce the Haar-Laplacian and discuss its properties, showing its advantages over existing Laplacians. We also present its properties for simple graphs. Section 4 is dedicated to the description of the graph learning problems that we tackle and to the presentation of the solution based on the Haar-Laplacian, the HaarNet network. Section 5 contains the numerical results obtained for learning (three types of link prediction problems) and signal processing (denoising). Section 6 concludes the paper.
@@ 2 Context
The core of our discussion is a key concept in spectral graph theory, namely the Graph Fourier Transform (GFT). The transformation itself leverages the eigenvectors of the Laplacian to perform signal processing on graphs. Essentially, the GFT enables the analysis and manipulation of signals defined over graph nodes, transforming them from the vertex domain to a frequency domain that reflects the underlying graph structure. This notion is employed both in the learning and signal processing matters that we will further discuss throughout this paper.
@@
where $U=\left[u_{0},...,u_{N-1}\right]$ is the matrix of eigenvectors (Fourier basis) corresponding to the diagonal matrix $\Lambda=\mathrm{diag}\left(\left[\lambda_{0},...,\lambda_{N-1}\right]\right)$ of eigenvalues (frequencies) of the Laplacian. As mentioned in the previous section, if $\mathcal{G}$ is undirected with nonnegative weights then it enjoys nice spectral properties. It has a full set of real and nonnegative eigenvalues with corresponding orthonormal eigenvectors, thus highlighting the meaning of $\Lambda$ and $U$ in the GFT definition.
@@
Using it in (3) or (4) gives symmetric positive semidefinite Laplacians. However, such a Laplacian catches only the average weight between two nodes; the differences are lost. It is in fact a brutal replacement of a directed graph with an undirected one, all information about orientation being lost. Such a Laplacian was used in [9], in the context of a spectral graph convolutional network (GCN), with good results in classification. When only connectivity matters, this Laplacian can be successfully used.
@@ 2.1 Learning perspective
Perhaps the most prominent association of the Laplacian within the frames of learning applications is with spectral graph convolutional networks (GCNs), which extend the convolutional neural networks (CNNs) to graph-structured data. Graphs lack an appropriate translation operation, therefore the convolution in the vertex domain is unfeasible. However, the convolution operation on graphs is enabled in the spectral domain by means of the graph Fourier transform (GFT) [10]. Then, in a manner similar to the classical signal processing, the convolution of two graph signals $x$ and $y$ in the spectral domain turns into a multiplication:
@@
where $U$ is the GFT matrix and $\odot$ is the element-wise (Hadamard) product.
This approach allows GCNs to extend the concept of spatial locality to the unique and irregular structure of graphs. In plain terms, a layer of a GCN is a convolution between a graph signal and a learnable filter $g_{\theta}$. A rudimentary $g_{\theta}$ could be a non-parametric filter, but such filter would not be localized [11]. Moreover, the convolution itself is a computationally intensive matrix multiplication and dealing with large graphs makes the Laplacian eigendecomposition a prohibitive operation. Hence, to address these issues, it is suggested to make use of a recursively-formulated polynomial evaluated in the Laplacian as a parametrization for the filter. The recurrence of Chebyshev polynomials can be exploited [12] to approximate $g_{\theta}$ as a truncated expansion, which further allows writing the filtering procedure as
@@
where $\theta_{k}$ are learnable filter coefficients and $\lambda_{max}$ is the largest eigenvalue of the normalized Laplacian $\bar{L}$. In [9], the simplification to a model that performs first-order local smoothing is proposed, which is computationally efficient and sufficient for many tasks. Additionally, $\lambda_{max}=2$ is employed as an approximation, this way eliminating the need to actually compute the largest eigenvalue.
@@
A drawback of the magnetic Laplacian is its lack of robustness to scaling. When multiplying $A$ with a constant $c$, the real part of (10) scales correctly; however, the phase is multiplied with $c$, so it may go outside the interval $[-\pi,\pi]$ and thus change its meaning. For example, take $a_{uv}=1$, $a_{vu}=0$ and so $h_{uv}=0.5(\cos(\pi/2)+i\sin(\pi/2))=0.5i$. Then, for the matrix $4A$ we would obtain $h_{uv}=2(\cos(2\pi)+i\sin(2\pi))=2$; this value is the same as that for an adjacency matrix with $a_{uv}=a_{vu}=2$, which has two antiparallel edges and thus could be associated with symmetry. Similarly, for the matrix $3A$, it results that $h_{uv}=1.5(\cos(3\pi/2)+i\sin(3\pi/2))=-1.5i$; hence, one may say that the edge is reversed (or the sign of the weight is flipped).
So, working with the magnetic Laplacian implies special care in choosing the value of the parameter $q$. For example, in [14], the value $q=1/2\max(a_{uv}-a_{vu})$ is taken, ensuring non-ambiguity of phase. In [15], the value of $q$ is chosen by appealing to Johnson’s algorithm for detecting directed cycles.
@@
However, while the sign-magnetic Laplacian captures perfectly the orientation of the edges, it is not devoid of criticism. Essentially, we would like that different graphs have different Laplacian, a fact that is true for undirected graphs; however, this is not true for the Laplacians in this section. In particular, the sign-magnetic Laplacian fails when both edges $(u,v)$ and $(v,u)$ exist. It is easy to see in (14) that the same value $h_{uv}$ is obtained by taking, for example, $a_{uv}=10000$ and $a_{vu}=1$ in one case and $a_{uv}=5001$ and $a_{vu}=5000$ in the other; in both cases, it results that $h_{uv}=5000.5i$; so, there is no sensitivity to variations that may be huge. Also, the sign-magnetic Laplacian is not continuous; if $a_{uv}=1$ and $a_{vu}=1$, then $h_{uv}=1$; if $a_{uv}=1.01$ and $a_{vu}=0.99$, then $h_{uv}=i$; there is a palpable discontinuity even though the change of weights is minor. Finally, the sign-magnetic Laplacian is zero when $A$ is skew-symmetric, since $A_{s}=0$.
We thus appreciate that the sign-magnetic Laplacian is not necessarily the best option for graphs with pairs of antiparallel edges, like those modeling money transfers (it matters how imbalanced are the transfers between two accounts, not only the average transferred amount) or trust networks; if trust is evaluated on a scale from $-10$ to $10$, having $a_{uv}=10$ and $a_{vu}=-10$ means that $u$ totally trusts $v$, but $v$ has no confidence at all in $u$; this can be a stab-in-the-back situation; while $a_{uv}=1$ and $a_{vu}=-1$ might mean that $u$ and $v$ barely know each other; however, both cases are coded with the same value, 00, in the sign-magnetic Laplacian (as well as in the magnetic Laplacian).
@@
Another track of research, that we do not follow here, is based on Laplacians that are symmetric although the graph is directed. An example, based on a personalized PageRank approach, is [22].
@@ 3 The Haar-Laplacian
be the skew-symmetric matrix naturally generated from $A$.
@@
$D_{h}=\mathrm{diag}(|H|\cdot\bm{1}).$ (17)
@@
due to (17). Since the diagonal elements of the matrix are zero, Gershgorin’s circle theorem [33] says that its eigenvalues lie inside the unit disk. It results immediately that the eigenvalues of (19) are in $[0,2]$ and hence the matrix (19) is positive semidefinite. Since
@@
The properties discussed critically in Section 2.1 for other Laplacians hold trivially for the Haar-Laplacian (18). Some of the properties follow from the fact that no information is lost from $A$ to $H_{h}$; the map $A\longrightarrow H_{h}$ is one-to-one; none of the previous Laplacians has this property.
@@
When the matrix $A$ is symmetric, $H_{h}=A_{s}$ is real and symmetric and its eigenvectors are real. The Haar-Laplacian is identical with the standard Laplacian (3) (and so are the magnetic Laplacians).
@@
$L_{h}=U\Lambda U^{H}$ (21)
be the eigenvalue decomposition of the Haar-Laplacian (18), where $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$, $\lambda_{1}\leq\ldots\leq\lambda_{n}$, and $U\in{\mathbb{C}}^{n\times n}$ is unitary and contains on columns the eigenvectors $u_{1},\ldots,u_{n}$. In GSP context, the matrix $U$ is called GFT and the eigenvalues $\lambda_{j}\geq 0$ are named graph frequencies. As discussed in [23] and other works, smaller eigenvalues are associated with lower frequencies.
@@
The corresponding Haar-Laplacian is
$L_{h}=\sqrt{2}I-\frac{1}{2}(S+S^{T})-i\frac{1}{2}(S-S^{T})=\sqrt{2}I-\frac{1+i}{2}S-\frac{1-i}{2}S^{T}.$
Like (22), it is also a circulant matrix, hence the GFT $U_{h}$ from (21) is the discrete Fourier transform matrix $W$. So, the Fourier transformation of this graph is the same as in classical signal processing, like for the random walk Laplacian [25]. However, the frequencies are different. Since $S$ is also diagonalized by $W$ and its eigenvalues are $\text{exp}(2\pi ij/n)$, $j=0:n-1$, the eigenvalues of $L_{h}$ are
$\sqrt{2}-\cos(2\pi j/n)-\sin(2\pi j/n),\ j=0:n-1.$ (23)
They take values in the interval $[0,\sqrt{2}]$. Due to the symmetries of the function $\cos\xi+\sin\xi$ with respect to $\pi/4$ and $45\pi/$ 4, the $n$ values (23) are distinct when $n$ is not a multiple of 4; otherwise, excepting the smallest and the largest, they appear in pairs.
For the adjacency matrix of a directed cycle graph, as well as for matrices that have a single value of 1 on each row, one can see from (10)-(12) and (18) that, if $q=1/8$, the Haar and magnetic Laplacians are related through $L_{h}=\sqrt{2}L_{m}$. So, the two Laplacians produce essentially identical frequencies. However, for other values of $q\leq 1/2$ they are different.
It is also immediate that for $q=1/4$, the magnetic and sign-magnetic Laplacians are identical for the directed cycle graph. For other values of $q\leq 1/2$, they are different.
Figure 1 shows the sorted frequencies of the Haar-Laplacian as well as those of the SVD [28], magnetic [6] and sign-magnetic [7] Laplacians for $n\in\{24,26\}$. The distribution of frequencies is similar, with a slight advantage for the Haar-Laplacian.
Figure 1: Left: sorted frequencies for Haar, SVD, magnetic ($q=1/4$) and sign-magnetic Laplacians for $n=24$ (the latter two curves are identical). Right: same, with $n=26$, $q=1/8$.
@@
Refer to caption Refer to caption Refer to caption
The sorted frequencies associated with the GFTs of the Haar, SVD, magnetic (with $q=1/8$) and sign-magnetic Laplacians are shown in Figure 3. There are insignificant differences between the three sets of frequencies; in particular, the magnetic and sign-magnetic Laplacians have the same eigenvalues for G15a. (Note that for $q=1/4$ the magnetic and sign-magnetic Laplacians are again identical.) The frequencies are nicely spread, so we can consider that all four Laplacians pass this sanity check.
Refer to caption Refer to caption Refer to caption

### v2

@@ Abstract
This paper introduces a novel Laplacian matrix aiming to enable the construction of spectral convolutional networks and to extend the signal processing applications for directed graphs. Our proposal is inspired by a Haar-like transformation and produces a Hermitian matrix which is not only in one-to-one relation with the adjacency matrix, preserving both direction and weight information, but also enjoys desirable additional properties like scaling robustness, sensitivity, continuity, and directionality. We take a theoretical standpoint and support the conformity of our approach with spectral graph theory. Then, we address two use cases: graph learning (by introducing HaarNet, a spectral graph convolutional network built with our Haar-Laplacian) and graph signal processing. We show that our approach gives better results in applications like weight prediction and denoising on directed graphs.
@@
graph Laplacian, directed graphs, graph neural networks, graph signal processing
@@ 1 Introduction
A multigraph can be accommodated to this framework by defining $w_{uv}$ as the sum of the weights of all edges between $u$ and $v$.
We assume that the graph does not have self-loops, hence $a_{uu}=0$, $\forall u\in\mathcal{V}$. We also assume that the graph is connected; there are no isolated vertices.
@@
where $|A|$ is the matrix whose elements are $|a_{uv}|$ and $\bm{1}$ is a vector with all elements equal to 1. When all weights are equal to one, this is the degree matrix: $d_{uu}$ is the number of edges containing vertex $u$.
The Laplacian associated with the graph is
L=DAL=D-A (3)
@@
If the graph is undirected with nonnegative weights, the Laplacian enjoys nice spectral properties, which proved useful in graph signal processing [1, 2, 3]. In particular, both Laplacians (3) and (4) are symmetric and positive semidefinite; the eigenvalues of the normalized Laplacian (4) lie in the interval $[0,2]$. The case of negative weights in undirected graphs is covered in [4].
Contribution. Here we propose a new construction, called Haar-Laplacian, based on the symmetrized and skew-symmetrized adjacency matrices associated with a directed graph. Like other Laplacians for directed graphs [5, 6, 7], the Haar-Laplacian is a Hermitian matrix and thus has real eigenvalues and orthogonal eigenvectors. We show that the Haar-Laplacian enjoys some properties that former Laplacians do not share. Moreover, we use the Haar-Laplacian in two types of applications. For graph learning, we introduce a spectral graph convolutional network called HaarNet and show its behavior for several instances of the link prediction problem in directed graphs; the best results are obtained for weight prediction, where the one-to-one relation of the Haar-Laplacian with the adjacency matrix gives it an edge over the existing approaches that do not have this property. We also illustrate the benefits of a variant of the Haar-Laplacian, with differences only in the diagonal, but with a sound frequency domain interpretation, in graph signal processing tasks such as denoising A related work is [8], where we have used the symmetrized and skew-symmetrized adjacency matrices for building separate blocks in an autoencoder network dedicated to anomaly detection; no connection with the Laplacian was made; no results, theoretical or experimental, are replicated in this paper.
Contents. In Sec. 2, we review the state-of-the-art in the topic of Laplacians for directed graphs, from two perspectives, those of graph learning and graph signal processing. In Sec. 3, we introduce the Haar-Laplacian and discuss its properties, showing its advantages over existing Laplacians. We also present its properties for simple graphs. Section 4 is dedicated to the description of the graph learning problems that we tackle and to the presentation of the solution based on the Haar-Laplacian, the HaarNet network. Section 5 contains the numerical results obtained for learning (three types of link prediction problems and a node classification problem) and signal processing (denoising). Section 6 concludes the paper.
@@ 2 Context
The core of our discussion is a key concept in spectral graph theory, namely the Graph Fourier Transform (GFT). The transformation itself leverages the eigenvectors of the Laplacian to perform signal processing on graphs. Essentially, the GFT enables the analysis and manipulation of signals defined over graph nodes, transforming them from the vertex domain to a frequency domain that reflects the underlying graph structure. This notion is employed in both the learning and signal processing matters that we will discuss in further detail throughout this paper.
@@
where $U=\left[u_{0},...,u_{N-1}\right]$ is the matrix of eigenvectors (Fourier basis) corresponding to the diagonal matrix $\Lambda=\mathrm{diag}\left(\left[\lambda_{0},...,\lambda_{N-1}\right]\right)$ of eigenvalues (frequencies) of the Laplacian. As mentioned in the previous section, if $\mathcal{G}$ is undirected with nonnegative weights, then it enjoys nice spectral properties. It has a full set of real and nonnegative eigenvalues with corresponding orthonormal eigenvectors, thus highlighting the meaning of $\Lambda$ and $U$ in the GFT definition.
@@
Using it in (3) or (4) gives symmetric positive semidefinite Laplacians. However, such a Laplacian only captures the average weight between two nodes; the differences are lost. It is in fact a brutal replacement of a directed graph with an undirected one, all information about orientation being lost. Such a Laplacian was used in [9], in the context of a spectral graph convolutional network (GCN), with good results in classification. When only connectivity matters, this Laplacian can be successfully used.
@@ 2.1 Learning perspective
Perhaps the most prominent association of the Laplacian within the frames of learning applications is with spectral graph convolutional networks (GCNs), which extend the convolutional neural networks (CNNs) to graph-structured data. Graphs lack an appropriate translation operation; therefore, the convolution in the vertex domain is unfeasible. However, the convolution operation on graphs is enabled in the spectral domain by means of the graph Fourier transform (GFT) [10]. Then, in a manner similar to the classical signal processing, the convolution of two graph signals $x$ and $y$ in the spectral domain turns into a multiplication:
@@
where $U$ is the GFT matrix in (5) and $\odot$ is the element-wise (Hadamard) product.
This approach allows GCNs to extend the concept of spatial locality to the unique and irregular structure of graphs. In plain terms, a layer of a GCN is a convolution between a graph signal and a learnable filter $g_{\theta}$. A rudimentary $g_{\theta}$ could be a nonparametric filter, but such a filter would not be localized [11]. Moreover, the convolution itself is a computationally intensive matrix multiplication, and dealing with large graphs makes the Laplacian eigendecomposition a prohibitive operation. Hence, to address these issues, it was suggested to make use of a recursively-formulated polynomial evaluated in the Laplacian as a parametrization for the filter. The recurrence of Chebyshev polynomials can be exploited [12] to approximate $g_{\theta}$ as a truncated expansion, which further allows writing the filtering procedure as
@@
where $\theta_{k}$ are learnable filter coefficients and $\lambda_{max}$ is the largest eigenvalue of the normalized Laplacian $\bar{L}$ (4). In [9], the simplification to a model that performs first-order local smoothing is proposed, which is computationally efficient and sufficient for many tasks. Additionally, $\lambda_{max}=2$ is employed as an approximation, this way eliminating the need to actually compute the largest eigenvalue.
@@
A drawback of the magnetic Laplacian is its lack of robustness to scaling. When multiplying $A$ with a constant $c$, the real part of (10) scales correctly; however, the phase is multiplied by $c$, so it may go outside the interval $\pi,\pi]$ and thus change its meaning. For example, take $a_{uv}=1$, $a_{vu}=0$ and so $h_{uv}=0.5(\cos(\pi/2)+i\sin(\pi/2))=0.5i$. Then, for the matrix $4A$ we would obtain $h_{uv}=2(\cos(2\pi)+i\sin(2\pi))=2$; this value is the same as that of an adjacency matrix with $a_{uv}=a_{vu}=2$, which has two antiparallel edges and could thus be associated with symmetry. Similarly, for the matrix $3A$, it results that $h_{uv}=1.5(\cos(3\pi/2)+i\sin(3\pi/2))=-1.5i$; hence, one may say that the edge is reversed (or the sign of the weight is flipped).
So, working with the magnetic Laplacian implies special care in choosing the value of the parameter $q$. For example, in [14], where the magnetic Laplacian is generalized for negative weights, the value $q=1/2\max|a_{uv}-a_{vu}|$ is taken, ensuring the non-ambiguity of the phase. In [15], the value of $q$ is chosen by appealing to Johnson’s algorithm for detecting directed cycles. However, these methods are efficient when data properties are known; the behavior for new data may not be robust.
@@
However, while the sign-magnetic Laplacian captures perfectly the orientation of the edges, it is not devoid of criticism. Essentially, we would like that different graphs have different Laplacians, a fact that is true for undirected graphs; however, this is not true for the Laplacians in this section. In particular, the sign-magnetic Laplacian fails when both edges $(u,v)$ and $(v,u)$ exist. It is easy to see in (14) that the same value $h_{uv}$ is obtained by taking, for example, $a_{uv}=10000$ and $a_{vu}=1$ in one case and $a_{uv}=5001$ and $a_{vu}=5000$ in the other; in both cases, it results that $h_{uv}=5000.5i$; so, there is no sensitivity to variations that may be huge. Also, the sign-magnetic Laplacian is not continuous; if $a_{uv}=1$ and $a_{vu}=1$, then $h_{uv}=1$; if $a_{uv}=1+\epsilon$ and $a_{vu}=1-\epsilon$, with $\epsilon>0$ then $h_{uv}=i$; there is a palpable discontinuity even though the change of weights may be minor. Finally, the sign-magnetic Laplacian is zero when $A$ is skew-symmetric, since $A_{s}=0$.
We thus appreciate that the sign-magnetic Laplacian is not necessarily the best option for graphs with pairs of antiparallel edges, like those modeling money transfers (it matters how imbalanced are the transfers between two accounts, not only the average transferred amount) or trust networks; if trust is evaluated on a scale from 10-10 to $10$, having $a_{uv}=10$ and $a_{vu}=-10$ means that $u$ totally trusts $v$, but $v$ has no confidence at all in $u$; this can be a stab-in-the-back situation; while $a_{uv}=1$ and $a_{vu}=-1$ might mean that $u$ and $v$ barely know each other; however, both cases are coded with the same value, 0, in the sign-magnetic Laplacian (as well as in the magnetic Laplacian).
@@
Another track of research, that we do not follow here, is based on Laplacians that are symmetric although the graph is directed. An example, based on a personalized PageRank approach, is [22]. Also aside our interests are approaches that only approximate the Laplacian [23].
@@ 3.1 Definition and properties
be the skew-symmetric matrix naturally generated from the adjacency matrix $A$.
@@
$D_{h}=\mathrm{diag}(|H_{h}|\cdot\bm{1}).$ (17)
@@
due to (17). Since the diagonal elements of the matrix $\bar{H}_{h}$ are zero, Gershgorin’s circle theorem [34] says that its eigenvalues lie inside the unit disk. It results immediately that the eigenvalues of (19) are in $[0,2]$ and hence the matrix (19) is positive semidefinite. Since
@@
The properties discussed critically in Section 2.1 for other Laplacians hold trivially for the Haar-Laplacian (18). Some of the properties follow from the fact that no information is lost from $A$ to $H_{h}$; the map $A\longrightarrow H_{h}$ is one-to-one; none of the previous Laplacians has this property. (Note that self-loops would allow several graphs to have the same Laplacians, which would be detrimental in learning problems.)
@@
When the matrix $A$ is symmetric, $H_{h}=A_{s}$ is real and symmetric and its eigenvectors are real. The Haar-Laplacian (18) is identical with the standard Laplacian (3) (and so is the magnetic Laplacian).
@@
3.2 Frequency domain interpretation
Let $z\in{\mathbb{C}}^{N}$ be a signal on the graph. For an undirected graph with positive weights (we assume $a_{uv}\geq 0$ for the discussion below), a measure of the variability of a signal on the graph is the total variation (TV)
$\text{TV}_{2}(z)=\frac{1}{2}\sum_{u,v=1}^{N}a_{uv}d(z_{u},z_{v})=z^{H}Lz,$ (21)
where $d(z_{u},z_{v})=\left|z_{u}-z_{v}\right|^{2}$ and $L$ is the Laplacian (3). In GSP context, the eigenvalues of $L$ are named graph frequencies and the matrix of eigenvectors is called GFT. As discussed in [24] and other works, smaller eigenvalues are associated with lower frequencies, since obviously TV is small in this case. In particular, since $\bm{1}$ is the eigenvector corresponding to the eigenvalue 0, it results that $\text{TV}_{2}(\bm{1})=0$. Another way of expressing the squared difference between two signal values is
$d(z_{u},z_{v})=\left[\begin{array}[]{cc}z_{u}^{*}&z_{v}^{*}\end{array}\right]Q\left[\begin{array}[]{c}z_{u}\\ z_{v}\end{array}\right].$ (22)
Relation (21) is given by $Q=\left[\begin{array}[]{cc}1&-1\\ -1&1\end{array}\right]$. Note the $Q$ is positive semidefinite; one of the eigenvalues is 0 and the corresponding eigenvector is $\bm{1}$.
When the graph is directed, TV (21) takes into account only the symmetric part of the adjacency matrix $A$. To involve the antisymmetric part and thus make a connection with Hermitian Laplacians, we can redefine (22) by using a complex matrix
$Q=\left[\begin{array}[]{cc}1&\alpha-i\beta\\ \alpha+i\beta&1\end{array}\right],$ (23)
where $\alpha,\beta\in{\mathbb{R}}$ are parameters that can be chosen. Since we want TV to be zero for a constant vector, the condition $0=\bm{1}^{T}Q\bm{1}=2+2\alpha$ gives $\alpha=-1$. For such $\alpha$, the matrix $Q$ is no longer positive semidefinite if $\beta\neq 0$. However, we can define a TV measure by using the absolute value of (21):
$\text{TV}_{Q}(z)=\left|\frac{1}{2}\sum_{u,v=1}^{N}a_{uv}d(z_{u},z_{v})\right|,$ (24)
together with (22) and (23). We can now make the connection with a slightly modified Haar-Laplacian.
Proposition 9
For a graph with positive weights, let us define the Laplacian
$\tilde{L}_{h}=D_{s}-H_{h},$ (25)
where $D_{s}$ is defined in (9) and $H_{h}$ in (16). Note that this is the Haar-Laplacian with a different diagonal; we will call (25) the HaarD-Laplacian. Then, for all $z\in{\mathbb{C}}^{N}$, the following equality holds:
$\text{TV}_{Q}(z)=\left|z^{H}\tilde{L}_{h}z\right|,$ (26)
where $\text{TV}_{Q}$ is defined in (24), with $Q$ as in (23) and $\alpha=-1$, $\beta=1$.
Proof. Relation (26) can be proved directly, by expliciting the two sides. Details are shown in Appendix A.
Remark 10
@@ 3.2 Frequency domain interpretation
$L_{h}=U\Lambda U^{H},\ \tilde{L}_{h}=\tilde{U}\tilde{\Lambda}\tilde{U}^{H}$ (27)
be the eigenvalue decompositions of the Haar- (18) and HaarD-Laplacian (25), where $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{N})$ and $U\in{\mathbb{C}}^{N\times N}$ is unitary and contains on columns the eigenvectors $u_{1},\ldots,u_{N}$; $\tilde{\Lambda}$ and $\tilde{U}$ are defined similarly.
The relation (26) between total variation and the HaarD-Laplacian suggests that, seen as graph frequencies, the eigenvalues of $\tilde{\Lambda}$ should be ordered in increasing order of their absolute value and numbered such that $|\lambda_{1}|\leq\ldots\leq|\lambda_{N}|$; remind that they are not necessarily positive. Lacking any other information, the eigenvalues of the Haar-Laplacian are numbered increasingly: $\lambda_{1}\leq\ldots\leq\lambda_{N}$; they are all nonnegative.
Note that although $\bm{1}^{T}\tilde{L}_{h}\bm{1}=0$ (since $\text{TV}_{Q}(\bm{1})=0$ by the choice of the matrix $Q$ in (23)), the vector $\bm{1}$ is not necessarily an eigenvector of the HaarD-Laplacian. On the downside, in general, there are other vectors $z$ for which $\text{TV}_{Q}(z)=0$, although their elements are not equal. So, the interpretation of low frequencies may be somewhat fuzzy. However, at the other end of the spectrum, large frequencies are clearly related with a high value of $\text{TV}_{Q}$ (24).
Remark 11
For a real signal $z\in{\mathbb{R}}^{N}$, relation (26) becomes
$\text{TV}_{Q}(z)=z^{T}(D_{s}-A_{s})z\geq 0,$ (28)
with $D_{s}$ from (9) and $A_{s}$ from (6). This is due to the fact that $z^{T}A_{a}z=0$ for a skew-symmetric matrix $A_{a}$. The matrix $D_{s}-A_{s}$ is the Laplacian associated with the symmetrized adjacency matrix and is positive semidefinite. This is another way of saying that $\text{TV}_{Q}$ reduces to $\text{TV}_{2}$ for real signals.
@@
The corresponding Haar- and HaarD-Laplacians have the form
$\gamma I-\frac{1+i}{2}S-\frac{1-i}{2}S^{T},$ (30)
with $\gamma=\sqrt{2}$ for Haar and $\gamma=1$ for HaarD. Like (29), the matrix (30) is also circulant, hence the GFTs $U_{h}$ and $\tilde{U}_{h}$ in (27) are permutations of the discrete Fourier transform matrix $W$. So, the Fourier transform of this graph is basically the same as in classical signal processing, like for the random walk Laplacian [26]. However, the frequencies are different. Since $S$ in (29) is also diagonalized by $W$ and its eigenvalues are $\text{exp}(2\pi ij/n)$, j=0:N1j=0:N-1, the eigenvalues of the Laplacians (30) are
$\gamma-\cos(2\pi j/N)-\sin(2\pi j/N),\ j=0:N-1.$ (31)
They take values in the interval $\gamma-\sqrt{2},\gamma+\sqrt{2}]$. Due to the symmetries of the function $\cos\xi+\sin\xi$ with respect to $\pi/4$ and $5\pi/4$, the $N$ values (31) are distinct when $N$ is not a multiple of 4; otherwise, excepting the smallest and the largest, they appear in pairs.
Figure 1 (top) shows the sorted frequencies (see Remark 10) of the Haar, HaarD and magnetic Laplacians (for $q=1/4$, one of the most popular choices) for $N=24$. The distribution of frequencies is similar. Note that for $q=1/4$, the magnetic [6] and sign-magnetic [7] Laplacians are identical for the directed cycle graph. For other values of $q\leq 1/2$, they are generally different.
Figure 1 (bottom) shows the TV (24) for the eigenvectors of the above Laplacians. For HaarD, the curve is identical with that from the top figure, since $\text{TV}_{Q}(\tilde{u}_{j})=|\tilde{\lambda}_{j}|$, see (26), and the eigenvalues are sorted in increasing order of absolute value. For the Haar-Laplacian, the order is different; the low frequencies correspond to low TV values, but not in increasing order; however, the high frequency eigenvectors are the same as for the HaarD-Laplacian. The TV for the magnetic Laplacian eigenvectors is generally increasing as eigenvalues increase, but not monotonically; for other values of $N$ the behavior can be even more chaotic.
Figure 1: Top: sorted frequencies for Haar, HaarD, and magnetic ($q=1/4$) Laplacians for $N=24$. Bottom: total variation (24) of the eigenvectors of the same Laplacians.